OpenStudy (anonymous):

Tanya is training a turtle for a turtle race. For every of an hour that the turtle is crawling, he can travel of a mile. At what unit rate is the turtle crawling?

OpenStudy (danjs):

Rate = distance / time

OpenStudy (anonymous):

For every 1/3 of an hour he can travel 1/12 of a mile.

OpenStudy (danjs):

divide the distance over time, to get a distance per 1 unit of time

OpenStudy (danjs):

\[\frac{ 1/12 mi }{ 1/3 hr }\]

OpenStudy (danjs):

= \[[\frac{ 1 }{ 12 }\frac{ 3 }{ 1 }]\frac{ mi }{ hr }\]

OpenStudy (anonymous):

I know how to do it. I just want to get my answered checked

OpenStudy (anonymous):

Is my answer going to be a whole number or a fraction?

OpenStudy (anonymous):

I got a fraction as my answer

OpenStudy (danjs):

1/4 mile / 1 hr

OpenStudy (anonymous):

Yeah, 1/4 of a mile per hour. So its right. Thanks!

OpenStudy (danjs):

just thinking about it, in 1/3 hr he goes 1/12 mi, so in 3 times that time, or 1 hr (3/3) he goes 3 times 1/12 miles, or 3/12 (1/4

OpenStudy (anonymous):

I can just submit 1/4 and let you know if it's right. I don't think it would be 3 time though. It could be 3/12

OpenStudy (anonymous):

Its right 1/4. I just submitted it. Thanks for the help!

OpenStudy (danjs):

1/4 is right i was just saying if he goes 1/12 in 1/3 hrs, in 3 times that time he goes 3/12 mi in 3/3 hr , OR 1/4 miles in 1 hr

OpenStudy (anonymous):

I just became your fan.-I think :/ LOL

OpenStudy (danjs):

That is just thinking about it, mathematically you just divide the distance by the time, to get a distance per unit of 1 time. (1/12)/(1/3) = 1/4

OpenStudy (anonymous):

Yep :)